solve by the substatution method
7x-8y=-35
7x+84=y

Answer by Alan3354(69443)

7x-8y=-35
7x+84=y --> 7x = y-84
Sub for 7x in the 1st eqn
y-84 - 8y = -35
-7y = -119
y = 17
----------
Sub for y and find x